# Fractal Patterns in Discrete Laplacians: Iterative Construction on 2D Square Lattices

### v1

@@ Abstract
We investigate the iterative construction of discrete Laplacians on 2D square lattices, revealing emergent fractal-like patterns shaped by modular arithmetic. While classical $2$-style iterations reproduce known structures such as the Sierpiński triangle, our alternating binary–ternary (2322-style) process produces a novel class of aperiodic figures. These display low density variance, minimal connectivity loss, and non-repetitive organization reminiscent of Dekking’s sequences. Fourier and autocorrelation analyses confirm their quasi-periodic nature, suggesting applications in self-assembly, sensor networks, and biological modeling. The findings open new paths toward structured randomness and fractal dynamics in discrete systems.
These findings also open avenues for exploring higher-dimensional Laplacian constructions and their implications in quasicrystals, aperiodic tilings, and stochastic processes.
@@ Introduction
Previous studies of Discrete Laplacians on 2D lattices have documented visually striking fractal patterns, such as snowflake-like and carpet-like motifs, arising from binary iterations alone Aiba et al. (2006). These patterns emerge due to modular arithmetic constraints and exhibit diverse forms, depending on the chosen seed and neighborhood structure. However, binary (modulo 2) iterations tend to produce periodic structures and, at specific iterations, dissociate into their initial seed configurations. This motivates the exploration of higher-order modular arithmetic, such as ternary (modulo 3) and quaternary (modulo 4) iterations, which significantly expand the space of possible patterns and often result in aperiodic, structurally richer formations.
In this paper, we investigate a special class of iterative Laplacian constructions, focusing on the interplay between binary and ternary arithmetic in pattern formation. We introduce the 2322-style pattern, an alternating sequence of binary and ternary iterations, which gives rise to structurally distinct, non-repetitive figures. Unlike purely binary $2$-style figures, which exhibit dissociative periodicity, 2322-style figures display low density variance, minimal connectivity loss, and quasi-aperiodicity. These properties are reminiscent of Dekking’s non-repetitive sequences, which play a significant role in combinatorial mathematics and aperiodic order.
Beyond their mathematical significance, these structures may have applications in physics, materials science, and computational modeling. The concluding section of this paper will discuss their potential relevance, including connections to biological self-organization, sensor technology, material science, and fractal-based computational algorithms.
@@ 1.2 Automaton Evolution
The iterative update rule follows the form:
$\nu_{i}(p)=\sum_{g\in Ne(p)}\big{(}\nu(g)-\nu(p)\big{)}\mod n,$ (1)
where:
@@
    $n$ is the modulus (defining binary, ternary, or higher-order arithmetic),
@@
    $i$ represents the iteration step,
@@
    $Ne(p)$ defines the neighborhood structure, governing local interactions.
@@ 1.3 Modular Arithmetic and Iterative Extensions
To extend beyond binary figures, we introduce modular arithmetic with $n>2$. Each color represents a distinct modular class:
@@
    Binary iteration ($n=2$): classical black-and-white figures.
@@
    Ternary iteration ($n=3$): introduces a new state, increasing complexity.
  • 3.
    Quaternary and higher ($n>3$): potential multi-state structures.
We define a $k$-nary iteration:
$u_{i}(p)=\sum_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod k,\quad k=2,3,4,\dots$
where the iteration is binary for $k=2$, ternary for $k=3$, and so forth.
In this paper, we examine specific iteration sequences:
$u_{i}(p)=\begin{cases}\sum\limits_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod n,&\text{if }i\equiv 2\pmod{4},\\ \sum\limits_{g\in Ne(p)}\big{(}u(g)-u(p)\big{)}\mod 2,&\text{otherwise}.\end{cases}$ (2)
This generates $222n$ 22-style figures, including:
@@
    $2$-style (purely binary evolution),
@@
    2322-style (introducing ternary interference in specific steps).
Figure 3: Binary iterations from a Sierpiński seed and Neumann neighborhood.
@@ 2 Binary style figures and their properties
Prior work on binary style figures by Suzuki and Maegaito uncovered visually striking configurations, including snowflake-like, butterfly-like, and Persian carpet-like motifs Aiba et al. (2006). Theoretical developments, including fixed-point theorems, periodicity results, and connections to binomial and trinomial sequences have been found Hadlich et al. (2011).
@@
All binary-style figures posses a special feature which is described in:
Proposition. Every binary-style figure of arbitrary seed and neighborhood at iterations $i=8k$, $k=1,2,3,\ldots$ consists of a spread of its seeds and is at least 13-steps away from connectedness.
@@
Generalization to Other Neighborhoods
Although this proof is carried out for Diag-Neumann, all observed figures for various seeds and neighborhoods exhibited the same behavior. Therefore, it is highly likely that all $222n$ 22-style figures for even $n$ behave similarly, all the figures are reduced to a spread of seeds at iterations $8k$.
@@ 2.1 Density Measure
$\rho(i)=\frac{\sum_{p\in\text{lattice}}\text{sgn}(f(p))}{(3+2i)^{2}}.$
@@ 2.2 Sierpinski-like Triangles
Figure 4: Sierpinski-like triangle constructions obtained as $2$-style figures at iterations $i=2^{k+1}-1$, where $k=0,1,2,\ldots$. The same neighborhood applied to different seeds produces structurally similar results, though the fractal dimension (via box-counting) varies depending on the seed.
Additionally, we compare the $2$-style and 2322-style figures generated at the same subsequence of iterations in Fig. 5. Notably, while $2$-style figures form clear, structured Sierpinski-like patterns, the 2322-style construction introduces greater structural complexity, leading to an apparently more chaotic formation.
@@
Figure 5: Comparison of $2$-style and 2322-style constructions from the same seeds and neighborhoods at the iteration sequence $i=2^{k+1}-1$, where $k=0,1,\ldots,5$. The 2322-style construction exhibits more irregularity and disorder compared to the structured Sierpinski-like growth of $2$-style figures.
@@ 3 Unique Properties of $2322$-Style Figures
The 2322-style construction, unlike purely binary ($2$) figures, introduces ternary interactions, resulting in distinctive structural and dynamical properties. These include:
@@
    Non-repetitive structural variation, distinct from the periodicity seen in $2$ figures.
At higher iterations, 2322-style figures retain complexity, while higher-order $222n$ 22-figures (for $n=5,7,9$) tend to exhibit more uniform growth patterns (Fig. 6).
@@
[…]
@@ 3.1 Connectivity and Structural Complexity
Figure 7: Comparison of $2$ and 2322 evolutions from the same seed. Unlike $2$ figures, 2322-style figures maintain structural connectivity, preventing dissociation into dispersed seeds at iteration 8.
[…]
@@ 3.2 Density Variability and Aperiodic Behavior
The density evolution of $2$ and 2322 figures over time is compared in Fig. 8. Notably, $2$ figures undergo periodic density reductions at iterations $8k$, $k=1,2,\dots$, corresponding to their spread-of-seeds phenomenon. This is absent in 2322-style figures, which exhibit a more consistent density profile.
@@
(a) $2$-style and 2322-style figures over iterations. Local minima at iterations $8k$ correspond to seed dissociation in $2$ figures, a behavior not observed in 2322 figures.
[…]
@@ 3.3 Connection to Aperiodic Sequences
4 Statistical Comparison of $2$ and 2322 Sequences
This section investigates the numerical properties of discrete Laplacian sequences generated by $2$-style and 2322-style constructions at a fixed lattice cell. The goal is to determine whether these structures exhibit long-range correlations, spectral properties, and fractal characteristics, providing potential links to Dekking’s Construction and other non-periodic sequences.
@@ 4.1 Statistical Characteristics
Table 1: Comparison of entropy, statistical measures, and fractal dimensions for $2$ and 2322 sequences over 500 iterations.
The higher entropy of the 2322 sequence indicates greater unpredictability compared to the periodic $2$ sequence. Similarly, the increased variance and higher fractal dimension suggest that 2322 structures exhibit greater structural complexity.
@@ 4.2 Fourier Spectral Analysis
Figures 9 and 10 show the Fourier transform of the sequences. The $2$ sequence exhibits sharp peaks, indicative of strong periodicity, whereas the 2322 sequence has a broad, distributed spectrum, suggesting quasi-periodicity or hierarchical self-organization.
Figure 9: Fourier spectrum of the $2$ sequence. Sharp peaks indicate strong periodicity.
@@
The absence of clear periodic peaks in the 2322 case supports the hypothesis that it resembles non-periodic substitution sequences, such as those found in Dekking’s Construction. This aligns with known fractal systems and aperiodic tilings, where self-similar structures emerge without strict repetition.
@@ 4.3 Autocorrelation Analysis
    The $2$ sequence exhibits clear periodic correlations, confirming its structured, repeating nature—akin to the Sierpiński triangle.
@@
Figure 11: Autocorrelation function of the $2$ sequence. Periodic correlations indicate a structured, repeating pattern.
@@ 4.4 Connection to Dekking’s Construction
Dekking’s Construction Dekking (1979) generates non-periodic sequences that avoid Abelian repetitions. The 2322 sequence, with its high entropy, irregular autocorrelation, and absence of Fourier periodicity, aligns closely with this framework.
@@
    $2$ figures behave as classical fractals, with periodic behavior and predictable density fluctuations.
@@
4.5 Implications for Fractal and Aperiodic Systems
The observed aperiodicity and fractal-like behavior in 2322 sequences could have applications in:
  • 1.
    Quasicrystals and Aperiodic Order: The lack of strict periodicity is reminiscent of quasicrystals and other self-organizing materials.
  • 2.
    Computational Mathematics: Potential use in graph algorithms, hierarchical clustering, and AI-driven optimizations.
  • 3.
    Physical and Biological Systems: Non-repetitive structures appear in self-assembled molecular networks, biological growth patterns, and stochastic neural models.
The higher fractal dimension suggests that 2322-style figures do not conform to classical self-similar fractal models, reinforcing the idea that they belong to a distinct mathematical category, governed by non-trivial modular arithmetic and aperiodic order.
@@ 4.6 Future Research Directions
5 Conclusions and Applications
This study has explored the iterative construction of discrete Laplacians on 2D square lattices, focusing on the emergent properties of 2322-style patterns. Unlike $2$-style figures, which tend to display periodic behavior and seed dissociation at specific iterations, 2322-style figures exhibit quasi-aperiodicity, minimal connectivity loss, and low density variance. The self-similar yet non-repetitive nature of these figures suggests deep connections to aperiodic sequences, fractal geometry, and material organization.
Our findings have implications across mathematics, physics, biology, and computational science, reinforcing the role of discrete Laplacians as a framework for understanding complex hierarchical structures.
@@ 5.1 Mathematical Implications
    Aperiodic Sequences and Non-Repetitive Growth: The structural complexity of 2322-style figures aligns with Dekking’s Construction, which generates strongly non-repetitive sequences in combinatorial mathematics Dekking (1979). The alternating binary-ternary structure prevents periodic breakdown, making these figures distinct from traditional cellular automata.
@@
    Fractal Properties: Box-counting methods estimate a fractal dimension of figures of approximately between 0.9 to 1.87, reinforcing the self-similar yet non-strictly repeating nature of these figures.
@@
    Higher-Dimensional Extensions: Extending the iterative Laplacian process to 3D lattices could uncover higher-rank algebraic structures and new combinatorial sequences relevant to discrete geometry.
5.2 Physical and Material Science Applications
@@
    Quasicrystals and Aperiodic Order: The quasi-periodicity of 2322-style figures is reminiscent of quasicrystalline structures, where long-range order exists without strict translational symmetry Gell-Mann (1994).
@@
    Energy Dissipation in Materials: Valianti (2015) noted that modular arithmetic-based structures govern energy minimization and stochastic relaxation in complex alloys Valianti et al. (2015). The structured, yet non-repetitive, growth of 2322 figures suggests a possible role in fracture mechanics and self-healing materials.
@@
    Fractal-Based Antennas, Sensors and Packages: Fractal-like structures have been successfully applied in sensor technology and electromagnetic wave manipulation Punjala and Makki (2009). The controlled irregularity of 2322 figures could inspire new designs for frequency-tuned fractal antennas and of properties of packages Zhang et al. (2021).
5.3 Biological and Computational Relevance

### v2

@@ Abstract
We investigate the iterative construction of discrete Laplacians on 2D square lattices, revealing emergent fractal-like patterns shaped by modular arithmetic. While classical 2222-style iterations reproduce known structures such as the Sierpiński triangle, our alternating binary–ternary (2322-style) process produces a novel class of aperiodic figures. These display low density variance, minimal connectivity loss, and non-repetitive organization reminiscent of Dekking’s sequences. Fourier and autocorrelation analyses confirm their quasi-periodic nature, suggesting structural analogies with self-assembly processes and fractal-based design.
The results highlight the structural richness of modular Laplacian dynamics and motivate further study of non-periodic discrete systems.
@@ Introduction
Previous studies of Discrete Laplacians on 2D lattices have documented visually striking fractal patterns, such as snowflake-like and carpet-like motifs, arising from binary iterations alone [1]. These patterns emerge due to modular arithmetic constraints and exhibit diverse forms, depending on the chosen seed and neighborhood structure. However, binary (modulo 2) iterations tend to produce periodic structures and, at specific iterations, dissociate into their initial seed configurations. This motivates the exploration of higher-order modular arithmetic, such as ternary (modulo 3) and quaternary (modulo 4) iterations, which significantly expand the space of possible patterns and often result in aperiodic, structurally richer formations.
Earlier investigations of modular Laplacian dynamics in the purely binary setting are presented in [5].
In this paper, we investigate a special class of iterative Laplacian constructions, focusing on the interplay between binary and ternary arithmetic in pattern formation. We introduce the 2322-style pattern, an alternating sequence of binary and ternary iterations, which gives rise to structurally distinct, non-repetitive figures. Unlike purely binary 2222-style figures, which exhibit dissociative periodicity, 2322-style figures display low density variance, minimal connectivity loss, and quasi-aperiodicity. These properties are reminiscent of Dekking’s non-repetitive sequences, which play a significant role in combinatorial mathematics and aperiodic order.
Beyond their mathematical significance, these structures may have broader structural implications. The concluding section of this paper will discuss their potential relevance, including connections to biological self-organization, sensor technology, material science, and fractal-based computational algorithms.
@@ 1.2 Automaton Evolution
Let $G=(V,E)$ denote the square lattice graph, where $V$ is the set of lattice points and $E$ is determined by the chosen neighborhood structure. For $p\in V$, we denote by
$Ne(p)=\{g\in V:(p,g)\in E\}$
the set of neighboring vertices of $p$.
Consider a sequence
$\mathcal{S}=(k_{1},k_{2},k_{3},\ldots),\qquad k_{i}\in\{2,3,4,\ldots\}.$
The discrete graph Laplacian applied to $\nu_{i}$ is defined by
$(\Delta\nu_{i})(p)=\sum_{g\in Ne(p)}\big(\nu_{i}(g)-\nu_{i}(p)\big)$ (1)
defined over $\mathbb{Z}$ prior to modular reduction. The evolution rule is given by
$\nu_{i+1}=(\Delta\nu_{i})\mod k_{i},$
where $k_{i}$ is the $i$-th element of $\mathcal{S}$.
Here:
@@
    $k_{i}$ is the modulus (binary for $k_{i}=2$, ternary for $k_{i}=3$, etc.),
@@
    $i$ denotes the discrete time (iteration) step,
@@
    $Ne(p)$ defines the neighborhood structure (e.g., von Neumann or diagonal).
Thus, the automaton evolves by repeated application of the graph Laplacian, followed by reduction modulo $k_{i}$, for $i=1,2,3,\ldots$.
@@ 1.3 Modular Arithmetic and Iterative Extensions
The introduction of the sequence $\mathcal{S}$ allows us to extend the construction beyond binary (black-and-white) figures, which correspond to the case when $\mathcal{S}$ is constant, i.e. $\mathcal{S}=(2,2,2,\ldots)$.
In particular, for constant sequences $\mathcal{S}=(n,n,n,\ldots)$ we obtain:
@@
    Ternary iteration ($n=3$): a three-state structure with increased combinatorial complexity.
@@
    Quaternary and higher ($n>3$): multi-state structures with $n$ distinct states.
In this paper, we focus on sequences $\mathcal{S}$ of the form
$k_{i}=\begin{cases}n,&\text{if }i\equiv 2\pmod{4},\\ 2,&\text{otherwise},\end{cases}$ (2)
and the corresponding family of figures generated by the Laplacian dynamics will be referred to as $2n22$-style figures [^1] Other families of construction sequences are studied in [6].
In particular:
@@
    2222-style: purely binary evolution ($n=2$),
@@
    2322-style: ternary interference at every fourth step ($n=3$).
Figure 3: Binary ($2222$-style) evolution from a three-point seed under the von Neumann neighborhood. The upper left panel shows the initial seed and the assumed neighborhood structure. The subsequent panels display the 0-th (initial configuration), first, second, and third iterations generated by the modular Laplacian dynamics with $k_{i}=2$, illustrating the initial expansion induced by the discrete Laplacian.
@@ 2 Binary style figures and their properties
Prior work on binary style figures by Suzuki and Maegaito uncovered visually striking configurations, including snowflake-like, butterfly-like, and Persian carpet-like motifs [1]. Theoretical developments, including fixed-point theorems, periodicity results, and connections to binomial and trinomial sequences have been found [4].
@@
Proposition. Let the binary Laplacian dynamics be defined with the Diag-Neumann neighborhood and let the initial configuration be finitely supported. Then for every $k\geq 1$, the configuration at iteration $i=8k$ consists of disjoint translated copies of the initial seed. Moreover, these copies are separated by a positive gap (which equals $13$ in the $3\times 3$ case considered here).
@@
Remark on other neighborhoods.
The above argument is carried out for the Diag-Neumann neighborhood. Numerical experiments indicate that similar replication phenomena occur for other neighborhood structures as well. A general operator-based proof for arbitrary neighborhoods is beyond the scope of the present paper.
@@ 2.1 Density Measure
$\rho(i)=\frac{\sum_{p\in\text{lattice}}\text{sgn}(\nu(p))}{(3+2i)^{2}}.$
@@ 2.2 Sierpinski-like Triangles
Figure 4: Sierpinski-like triangle constructions obtained as 2222-style figures at iterations $i=2^{k+1}-1$, where $k=0,1,2,\ldots$. The same neighborhood applied to different seeds produces structurally similar results, though the fractal dimension (via box-counting) varies depending on the seed.
Additionally, we compare the 2222-style and 2322-style figures generated at the same subsequence of iterations in Fig. 5. Notably, while 2222-style figures form clear, structured Sierpinski-like patterns, the 2322-style construction introduces greater structural complexity, leading to an apparently more chaotic formation.
@@
Figure 5: Comparison of 2222-style and 2322-style constructions from the same seeds and neighborhoods at the iteration sequence $i=2^{k+1}-1$, where $k=0,1,\ldots,5$. The 2322-style construction exhibits more irregularity and disorder compared to the structured Sierpinski-like growth of $22$-style figures.
@@ 3 Unique Properties of $2322$-Style Figures
The 2322-style construction, unlike purely binary (2222) figures, introduces ternary interactions, resulting in distinctive structural and dynamical properties. These include:
@@
    Non-repetitive structural variation, distinct from the periodicity seen in 2222 figures.
At higher iterations, 2322-style figures retain complexity, while higher-order $2n22$-figures (for $n=5,7,9$) tend to exhibit more uniform growth patterns (Fig. 6).
@@
Figure 6: Seed: single point; diagonal Neumann neighborhood.
(a) Comparison of $2n22$ figures for $n=2,3,4,5$ (left to right; iterations 30–32,34). The 2322 case (second column) is structurally distinct. Fractal dimensions for 2222 (resp. 2322) are 1.31, 1.66, 0.91, 1.14 (resp. 1.78, 1.62, 1.43, 1.62).
(b) Comparison of 2322 with $2n22$ for odd $n=3,5,7,9$ (iterations 23–25,27), showing persistent structural distinctness.
@@ 3.1 Connectivity and Structural Complexity
Figure 7: Comparison of 2222 and 2322 evolutions from the same seed. Unlike 2222 figures, 2322-style figures maintain structural connectivity, in contrast to the seed dissociation observed in 2222 at 8ii-ths iterations.
Among 2nn22-style figures for $n$ odd, 2322 exhibits distinct structural behaviour:
@@ 3.2 Density Variability and Aperiodic Behavior
The density evolution of 2222 and 2322 figures over time is compared in Fig. 8. Notably, 2222 figures undergo periodic density reductions at iterations $8k$, $k=1,2,\dots$, corresponding to their spread-of-seeds phenomenon. This is absent in 2322-style figures, which exhibit a more consistent density profile.
@@
(a) 2222-style and 2322-style figures over iterations. Local minima at iterations $8k$ correspond to seed dissociation in 2222 figures, a behavior not observed in 2322 figures.
(b) 2nn22-style figures, $n=3,5,7,9$. Graphs for $n=5,7,9$ overlap, suggesting similar periodic behavior, whereas 2322 maintains a unique profile.
@@ 3.3 Connection to Aperiodic Sequences
4 Statistical Comparison of 2222 and 2322 Sequences
This section investigates the numerical properties of discrete Laplacian sequences generated by 2222-style and 2322-style constructions at a fixed lattice cell. The goal is to determine whether these structures exhibit long-range correlations, spectral properties, and fractal characteristics, providing potential links to Dekking’s Construction and other non-periodic sequences.
@@ 4.1 Statistical Characteristics
The entropy is computed in the Shannon sense,
$H=-\sum_{j}p_{j}\log p_{j},$
where $p_{j}$ denotes the empirical frequency of the $j$-th state. The fractal dimension is estimated using a standard box-counting method.
@@
Table 1: Comparison of entropy, statistical measures, and fractal dimensions for 2222 and 2322 sequences over 500 iterations.
The higher entropy of the 2322 sequence indicates greater unpredictability compared to the periodic 2222 sequence. Similarly, the increased variance and higher fractal dimension suggest that 2322 structures exhibit greater structural complexity.
@@ 4.2 Fourier Spectral Analysis
Figures 9 and 10 present the amplitude spectra of the corresponding sequences.
Figure 9: Fourier spectrum of the 2222 sequence. Sharp peaks indicate strong periodicity.
@@
The 2222 sequence exhibits pronounced discrete peaks located at regularly spaced frequencies, indicating strong periodic components and a well-defined fundamental frequency.
In contrast, the 2322 sequence does not display a single dominant periodic frequency. Instead, its spectrum consists of multiple distributed peaks with irregular spacing and varying amplitude, reflecting hierarchical spatial organization rather than strict periodic repetition.
Thus, while the 2222 case corresponds to a predominantly discrete spectrum, the 2322 sequence exhibits a more distributed spectral structure, consistent with non-periodic yet structured growth.
@@ 4.3 Autocorrelation Analysis
    The 2222 sequence exhibits clear periodic correlations, confirming its structured, repeating nature—akin to the Sierpiński triangle.
@@
Figure 11: Autocorrelation function of the 2222 sequence. Periodic correlations indicate a structured, repeating pattern.
@@ 4.4 Connection to Dekking’s Construction
Dekking’s Construction [2] generates non-periodic sequences that avoid Abelian repetitions. The 2322 sequence, with its high entropy, irregular autocorrelation, and absence of Fourier periodicity, shows structural similarities to this framework.
@@
    2222 figures behave as classical fractals, with periodic behavior and predictable density fluctuations.
@@ 4.5 Future Research Directions
5 Conclusions and Perspectives
We have studied modular Laplacian dynamics on two-dimensional square lattices, with particular emphasis on non-constant sequences such as the $2322$ pattern. In contrast to purely binary ($2222$) constructions, which exhibit periodic behavior and seed dissociation at specific iterations, the $2322$-style figures display reduced periodicity, higher entropy, and increased fractal dimension.
These results indicate that introducing variable modular arithmetic significantly modifies the long-term structural behavior of discrete Laplacian iterations.
@@ 5.1 Mathematical Implications
    Non-periodic growth: The structural complexity of $2322$-style figures suggests analogies with non-periodic substitution systems, including constructions related to Dekking-type sequences [2]. This correspondence is structural rather than formal.
@@
    Fractal characteristics: Box-counting estimates indicate fractal dimensions in the approximate range 0.90.91.871.87, consistent with the observed multi-scale self-similarity.
@@
    Higher-dimensional extensions: Extending modular Laplacian dynamics to three-dimensional lattices and other graph structures remains a natural direction for further study.
5.2 Physical and Computational Perspectives
@@
    Aperiodic order: The reduced periodicity of $2322$-style figures shows structural similarities to systems exhibiting aperiodic order, such as quasicrystalline patterns [3]. This analogy concerns geometric organization rather than a direct physical model.
@@
    Fractal-inspired design: Self-similar geometries are widely used in electromagnetic and sensor applications [7, 11]. The controlled irregularity of the $2322$-style figures suggests potential relevance for scalable pattern generation.
@@
    Structured discrete models: The non-uniform yet organized density patterns provide an abstract framework for studying hierarchical structures in graph-based systems.
  • 4.
    Motivation from natural self-assembly: Recent observations of Sierpiński-like protein structures in biological systems [8, 10, 9] demonstrate that fractal growth patterns may arise in natural self-assembly processes. While the present work is purely mathematical, such findings provide motivation for studying discrete self-similar dynamics.
5.3 Future Directions
Possible directions for further investigation include: